### v1

@@ 3.1 Complexity analysis
$\epsilon_{1}=\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{18}\epsilon^{1/2},$ (3.4)
in Lemma 2.4 and choosing $N_{t}$ in (2.8). For the case $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$, we have
$~{}\|g_{t+1}\|$ $\leqslant\frac{7}{3}\|g_{t}\|,$ (3.5)
$P(x_{t+1})$ $\leqslant P(x_{t})-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\bigg{(}\frac{-\sqrt{L_{2}\|g_{t}\|}+\sqrt{L_{2}\|g_{t}\|+72L_{2}\|g_{t+1}\|}}{6L_{2}}\bigg{)}^{2};$ (3.6)
@@
$P(x_{t+1})\leqslant P(x_{t})-\frac{7}{36\sqrt{L_{2}}}\|g_{t}\|^{3/2}.$ (3.7)
@@
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\frac{1}{2}\sqrt{L_{2}\|g_{t}\|})s_{t}-\frac{3}{4}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
@@
$\leqslant$ $~{}-\frac{5}{12}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{36}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-(\frac{7\sqrt{L_{2}}}{18}-\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\frac{\epsilon^{3/2}}{\sqrt{\|g_{t}\|}\|s_{t}\|})\sqrt{\|g_{t}\|}\|s_{t}\|^{2},$
$\leqslant$ $~{}-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}.$ (3.11)
@@
$\|g_{t+1}\|\leqslant\frac{1}{6}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|+\frac{L_{2}}{2}\|s_{t}\|^{2}.$ (3.12)
@@
$\|s_{t}\|\geqslant\frac{-\sqrt{L_{2}\|g_{t}\|}+\sqrt{L_{2}\|g_{t}\|+72L_{2}\|g_{t+1}\|}}{6L_{2}}.$ (3.13)
@@
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters chosen as in (3.4). For the case $\lambda_{\min}(H_{t})\leqslant-\dfrac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2},$ (3.14)
@@
$\|s_{t}\|=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}},~{}g_{t}^{\top}s_{t}\leqslant 0,~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\frac{1}{4\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2}.$ (3.16)
@@
$P(x_{t+1})-P(x_{t})\leqslant$ $-\frac{1}{12\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2}+\epsilon_{1}\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}+\frac{\epsilon_{2}}{2}\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}$
$\leqslant$ $-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|,\epsilon\}^{3/2},$
@@
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters choosen as in (3.4). For the case $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\epsilon}$, we have already found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$.
@@
$\bar{T}(\epsilon):=\min\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}\}.$
@@
$\mathcal{F}_{\bar{T}(\epsilon)}=\big{\{}t<\bar{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\big{\}},$ (3.17)
@@
From Lemmas 3.1 and 3.2, we conclude that each iteration $t$ belongs to one of the above two sets $\mathcal{F}_{\bar{T}(\epsilon)}$ and $\mathcal{G}_{\bar{T}(\epsilon)}$.
@@
$\epsilon_{1}=\min\{\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{18}\epsilon^{1/2},$ (3.19)
and choose $N_{t}$ in (2.8). We have
$T(\epsilon)\leqslant\lceil\log(\tilde{G}/\epsilon)/\log(2)\rceil|\mathcal{F}_{\bar{T}(\epsilon)}|+|\mathcal{F}_{\bar{T}(\epsilon)}|+1,$ (3.20)
where
$|\mathcal{F}_{\bar{T}(\epsilon)}|\leqslant 36\sqrt{L_{2}}(P(x_{0})-P^{*})\epsilon^{-3/2},$
$\tilde{G}:=\max\big{\{}\frac{7}{3}\tilde{C},1+\frac{L_{1}}{\sqrt{L_{2}}}\tilde{C}^{1/2}+\tilde{C},2+\frac{L_{1}}{\sqrt{L_{2}}}\big{\}}$ with $\tilde{C}:=\left(36\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}=\mathcal{O}(\rho^{1/3}\kappa)$, which means that $\tilde{G}=\mathcal{O}(\rho^{1/3}\kappa+\ell\rho^{-1/3})$.
@@
Remark 2
Theorem 3.1 shows that the iteration complexity of Algorithm 2 to find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\sqrt{L_{2}}\epsilon^{-1.5})=\tilde{\mathcal{O}}(\rho^{0.5}\kappa^{1.5}\epsilon^{-1.5})$. Moreover, we can also have a rough estimate of the number of inner iterations, i.e., $N_{t}$. If $x_{t}$ is not an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point, according to (3.2) and (3.9), we have $\|s_{t}\|\geqslant\mathcal{O}(\min\{\sqrt{\frac{\epsilon}{L_{2}}},\frac{\epsilon}{L_{1}+\sqrt{L_{2}}}\})$. By further combining with Lemma 2.4, we obtain that $A=\mathcal{O}(\rho^{0.5}\kappa^{0.5}\ell^{-2}\epsilon^{1.5}+\rho^{-0.5}\kappa^{-0.5}\epsilon)$, $N_{1}=\mathcal{O}(\kappa\log\frac{\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\kappa\log\frac{A+\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2)$. Following the same discussion as in Remark 1, through the Nesterov’s accelerated gradient algorithm, the convergence speed can be further improved to $N_{1}=\mathcal{O}(\sqrt{\kappa}\log\frac{\sqrt{\kappa+1}\|y_{0}-y^{*}(x_{0})\|}{A})$ and $N_{t}=\mathcal{O}(\sqrt{\kappa}\log(\sqrt{\kappa+1}+\frac{\sqrt{\kappa+1}\kappa\|s_{t-1}\|}{A})$ ($t\geqslant 2$).

### v2

@@ 3.1 Complexity analysis
$\epsilon_{1}=\min\bigg{\{}\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\bigg{\}}\epsilon^{3/2},~{}\epsilon_{2}=\frac{\sqrt{L_{2}}}{18}\epsilon^{1/2}$ (3.5)
in Lemma 2.4 and choosing $\eta_{y},\theta,N_{t}$ in (2.8). For the case $\|g_{t}\|\geqslant\epsilon$, if $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}$, we have
$~{}\|g_{t+1}\|\leqslant$ $~{}\frac{41}{9}\|g_{t}\|,$ (3.6)
$P(x_{t+1})\leqslant P(x_{t})-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\bigg{(}$ $\frac{-23\sqrt{L_{2}\|g_{t}\|}+\sqrt{529L_{2}\|g_{t}\|+648L_{2}\|g_{t+1}\|}}{18L_{2}}\bigg{)}^{2};$ (3.7)
@@
$P(x_{t+1})\leqslant P(x_{t})-\frac{5}{162\sqrt{L_{2}}}\|g_{t}\|^{3/2}.$ (3.8)
@@
$\leqslant$ $-\frac{1}{2}s_{t}^{\top}(H_{t}+\frac{1}{2}\sqrt{L_{2}\|g_{t}\|}I)s_{t}-\frac{3}{4}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
@@
$\leqslant$ $-\frac{5}{12}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\min\bigg{\{}\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\bigg{\}}\epsilon^{3/2}\|s_{t}\|+\frac{\sqrt{L_{2}}}{36}\epsilon^{1/2}\|s_{t}\|^{2}$
$\leqslant$ $-\left(\frac{7\sqrt{L_{2}}}{18}-\min\bigg{\{}\frac{1}{36},\frac{\sqrt{L_{2}}}{12L_{1}}\bigg{\}}\frac{\epsilon^{3/2}}{\sqrt{\|g_{t}\|}\|s_{t}\|}\right)\sqrt{\|g_{t}\|}\|s_{t}\|^{2},$
$\leqslant$ $-\dfrac{5}{18}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}.$ (3.12)
@@
$\|g_{t+1}\|\leqslant\frac{23}{18}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|+\frac{L_{2}}{2}\|s_{t}\|^{2}.$ (3.13)
@@
$\|s_{t}\|\geqslant\frac{-23\sqrt{L_{2}\|g_{t}\|}+\sqrt{529L_{2}\|g_{t}\|+648L_{2}\|g_{t+1}\|}}{18L_{2}}.$ (3.14)
@@
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters chosen the same as in Lemma 3.1. For the case $\lambda_{\min}(H_{t})\leqslant-\dfrac{1}{2}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\},$ (3.15)
@@
$\|s_{t}\|=\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}},~{}g_{t}^{\top}s_{t}\leqslant 0,~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}\leqslant-\frac{1}{4\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}.$ (3.17)
@@
$P(x_{t+1})\leqslant$ $~{}P(x_{t})-\frac{1}{12\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}+\epsilon_{1}\sqrt{\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}}+\frac{\epsilon_{2}}{2}\frac{\max\{\|g_{t}\|,\epsilon\}}{L_{2}}$
$\leqslant$ $~{}P(x_{t})-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\},$
@@
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters chosen the same as in Lemma 3.1. For the case $\|g_{t}\|\leqslant\epsilon$ and $\lambda_{\min}(H_{t})\geqslant-\dfrac{1}{2}\sqrt{L_{2}\epsilon}$, we have already found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$.
@@
$\bar{T}(\epsilon):=\min\left\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\lambda_{\min}(H_{t})\geqslant-\frac{1}{2}\sqrt{L_{2}\epsilon}\right\}.$
@@
$\mathcal{F}_{\bar{T}(\epsilon)}=\left\{t<\bar{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{36\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\right\},$ (3.18)
@@
From Lemmas 3.1 and 3.2, we conclude that each iteration $t$ belongs to at least one of the above two sets $\mathcal{F}_{\bar{T}(\epsilon)}$ and $\mathcal{G}_{\bar{T}(\epsilon)}$.
@@
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by Algorithm 2 with parameters chosen the same as in Lemma 3.1. We have
$T(\epsilon)$ $\leqslant\bar{C}_{1}\epsilon^{-3/2}\ln\epsilon^{-1}+\bar{C}_{2}\epsilon^{-3/2}+1=\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5}),$
$\sum_{t=0}^{T(\epsilon)}N_{t}\leqslant 2\sqrt{\kappa}$ $~{}\ln\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A}+2\sqrt{\kappa}T(\epsilon)\ln\frac{\sqrt{\kappa+1}(A+2\kappa\sqrt{\bar{G}}/\sqrt{L_{2}})}{A},$
where $\bar{C}_{1}=36\sqrt{L_{2}}(P(x_{0})-P^{*})/\ln 2$, $\bar{C}_{2}=(\ln\bar{G}+\ln 2)\bar{C}_{1}$, $\bar{G}:=\max\big{\{}\frac{41}{9}\bar{C},1+\frac{L_{1}}{\sqrt{L_{2}}}\bar{C}^{1/2}+\bar{C},2+\frac{L_{1}}{\sqrt{L_{2}}}\big{\}}$ with $\bar{C}:=\left(36\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}$, and $A$ is defined as in Lemma 2.4.
@@
Remark 3.1
Theorem 3.1 shows that the iteration complexity of Algorithm 2 to find an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point of $P(x)$ is $\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5})$. The total number of gradient ascent steps required is $\tilde{\mathcal{O}}(\ell^{2}\rho^{0.5}\mu^{-2}\epsilon^{-1.5})$. Compared to the GRTR algorithm, the LMNegCur algorithm has the same complexity results.
3.2 An inexact LMNegCur algorithm
In this subsection, we propose an inexact LMNegCur (ILMNegCur) algorithm that allows approximate solutions to the minimum eigenvalue subproblem (3.2) and the linear system subproblem (3.4) of algorithm 2 under certain criteria.
Specifically, to update $x$, the ILMNegCur algorithm applies the randomized Lanczos procedure Kuczynski1992Estimating to $H_{t}$, thereby obtaining the approximate minimum eigenvalue $\hat{\lambda}_{t}$ and unit vector $u_{t}$ of $H_{t}$, such that
$\hat{\lambda}_{t}\leqslant\lambda_{\min}(H_{t})+\frac{1}{4}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}},~{}\hat{\lambda}_{t}=u_{t}^{\top}H_{t}u_{t},~{}g_{t}^{\top}u_{t}\leqslant 0.$ (3.20)
If $\hat{\lambda}_{t}\leqslant-\frac{1}{4}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, then set
$s_{t}=\frac{1}{2}\sqrt{\max\{\|g_{t}\|,\epsilon\}/L_{2}}u_{t};$ (3.21)
Otherwise, if $\|g_{t}\|\geqslant\epsilon$ and $\hat{\lambda}_{t}\geqslant-\dfrac{1}{4}\sqrt{L_{2}\|g_{t}\|}$, then use the conjugate gradient method Royer2018ComplexityAO to solve the linear system subproblem (3.4) and get the direction $s_{t}$ such that
$\|g_{t}+(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)s_{t}\|\leqslant\frac{1}{4}\min\bigg{\{}\|g_{t}\|,\frac{1}{2}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|\bigg{\}}.$ (3.22)
Note that if $\|g_{t}\|\leqslant\epsilon$ and $\hat{\lambda}_{t}\geqslant-\dfrac{1}{4}\sqrt{L_{2}\epsilon}$, we have found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$. By replacing Steps 4 and 5 in the LMNegCur algorithm with the above steps, the detailed ILMNegCur algorithm can be immediately obtained.
3.3 Complexity analysis
Lemma 3.4
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by the ILMNegCur algorithm. Let
$\epsilon_{1}=\min\bigg{\{}\frac{1}{32},\frac{1}{144},\frac{\sqrt{L_{2}}}{32L_{1}}\bigg{\}}\epsilon^{3/2},~{}\epsilon_{2}=\min\bigg{\{}\frac{1}{12},\frac{1}{36}\bigg{\}}\sqrt{L_{2}}\epsilon^{1/2},$ (3.23)
and select $\eta_{y},\theta,N_{t}$ according to (2.4). In the $t$th iteration, if $\|g_{t}\|\geqslant\epsilon$ and $\hat{\lambda}_{t}\geqslant-\frac{1}{4}\sqrt{L_{2}\|g_{t}\|}$, then
$\|g_{t+1}\|\leqslant\frac{55}{8}\|g_{t}\|,$
$P(x_{t+1})\leqslant P(x_{t})-\dfrac{1}{24}\sqrt{L_{2}\|g_{t}\|}\bigg{(}\frac{-3\sqrt{L_{2}\|g_{t}\|}+\sqrt{9L_{2}\|g_{t}\|+8L_{2}\|g_{t+1}\|}}{2L_{2}}\bigg{)}^{2};$
If $\|g_{t+1}\|\geqslant\frac{1}{2}\|g_{t}\|$ holds, then we have
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{288\sqrt{L_{2}}}\|g_{t}\|^{3/2}.$ (3.24)
Proof
By a proof similar to (2.1), we have
$~{}P(x_{t+1})-P(x_{t})$
$\leqslant$ $~{}g_{t}^{\top}s_{t}+\frac{1}{2}s_{t}^{\top}H_{t}s_{t}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}$
$=$ $~{}s_{t}^{\top}(g_{t}+H_{t}s_{t}+\sqrt{L_{2}\|g_{t}\|}s_{t})-\frac{1}{2}s_{t}^{\top}(H_{t}+\frac{1}{2}\sqrt{L_{2}\|g_{t}\|}I)s_{t}-\frac{3}{4}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}$
$~{}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (3.25)
Combining (3.20) with $\hat{\lambda}_{t}\geqslant-\dfrac{1}{4}\sqrt{L_{2}\|g_{t}\|}$, we can easily see that
$\lambda_{\min}(H_{t})\geqslant\hat{\lambda}_{t}-\dfrac{1}{4}\sqrt{L_{2}\|g_{t}\|}\geqslant-\dfrac{1}{2}\sqrt{L_{2}\|g_{t}\|}.$ (3.26)
Then, by combining (3.3) with (3.22) and (3.26), we obtain that
$P(x_{t+1})-P(x_{t})\leqslant-\frac{5}{8}\sqrt{L_{2}\|g_{t}\|}\|s_{t}\|^{2}+\epsilon_{1}\|s_{t}\|+\frac{\epsilon_{2}}{2}\|s_{t}\|^{2}+\frac{L_{2}}{6}\|s_{t}\|^{3}.$ (3.27)
To further bound the right-hand side of (3.27), we first give some bounds on $\|s_{t}\|$. On the one hand, by combining (3.22) with (3.26), we then obtain that
$\|s_{t}\|=$ $~{}\|(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}(-g_{t}+g_{t}+(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)s_{t})\|$
$\leqslant$ $~{}\|(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)^{-1}\|\big{(}\|g_{t}\|+\frac{1}{4}\|g_{t}\|\big{)}\leqslant\frac{5}{2}\sqrt{\dfrac{\|g_{t}\|}{L_{2}}}.$ (3.28)
On the other hand, by further combining (3.22) with $\|H_{t}\|\leqslant\|H_{t}-\nabla^{2}P(x_{t})\|+\|\nabla^{2}P(x_{t})\|\leqslant L_{1}+\epsilon_{2}$, we obtain that
$\|s_{t}\|\geqslant\frac{\|g_{t}\|-\|g_{t}+(H_{t}+\sqrt{L_{2}\|g_{t}\|}I)s_{t}\|}{L_{1}+\epsilon_{2}+\sqrt{L_{2}\|g_{t}\|}}\geqslant\frac{\frac{3}{4}\|g_{t}\|}{L_{1}+\epsilon_{2}+\sqrt{L_{2}\|g_{t}\|}},$
which implies that
$\frac{4}{3}\epsilon_{2}\|s_{t}\|+\frac{4}{3}(L_{1}+\sqrt{L_{2}})\sqrt{\|g_{t}\|}\|s_{t}\|\geqslant\epsilon^{1/2}\|g_{t}\|\geqslant\epsilon^{3/2}.$ (3.29)
The rest of the proof is basically the same as Lemma 3.1. We omit it here for simplicity.
Lemma 3.5
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by the ILMNegCur algorithm with parameters chosen the same as in Lemma 3.4. At the $t$th iteration, if $\hat{\lambda}_{t}\leqslant-\frac{1}{4}\sqrt{L_{2}\max\{\|g_{t}\|,\epsilon\}}$, then
$P(x_{t+1})\leqslant P(x_{t})-\frac{1}{288\sqrt{L_{2}}}\max\{\epsilon^{3/2},\|g_{t}\|^{3/2}\},$ (3.30)
$\|g_{t+1}\|\leqslant 1+\frac{L_{1}}{2\sqrt{L_{2}}}\sqrt{\max\{\|g_{t}\|,\epsilon\}}+\max\{\|g_{t}\|,\epsilon\}.$ (3.31)
If $\|g_{t}\|\leqslant\epsilon$ and $\hat{\lambda}_{t}\geqslant-\dfrac{1}{4}\sqrt{L_{2}\epsilon}$, we have found an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point $x_{t}$.
The proof of Lemma 3.5 is basically the same as the proofs of Lemma 3.2 and Lemma 3.3, and will not be repeated here.
We are now ready to establish the iteration complexity for the ILMNegCur algorithm. In particular, let $\epsilon>0$ be any given target accuracy. We denote the first iteration index of the inexact LMNegCur algorithm to achieve $\|g_{t}\|\leqslant\epsilon$ and $\hat{\lambda}_{t}\geqslant-\frac{1}{4}\sqrt{L_{2}\epsilon}$ by
$\tilde{T}(\epsilon):=\min\left\{t|\|g_{t}\|\leqslant\epsilon~{}\text{and}~{}\hat{\lambda}_{t}\geqslant-\frac{1}{4}\sqrt{L_{2}\epsilon}\right\}.$
Then, by Lemma 3.5, we know that $x_{\tilde{T}(\epsilon)}$ is already an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point. To facilitate complexity analysis, we define the following indicator set:
$\mathcal{F}_{\tilde{T}(\epsilon)}=\left\{t<\tilde{T}(\epsilon):P(x_{t+1})-P(x_{t})\leqslant-\frac{1}{288\sqrt{L_{2}}}\max\{\|g_{t}\|^{3/2},\epsilon^{3/2}\}\right\},$ (3.32)
$\mathcal{G}_{\tilde{T}(\epsilon)}=\left\{t<\tilde{T}(\epsilon):\|g_{t+1}\|\leqslant\frac{1}{2}\|g_{t}\|,t\notin\mathcal{F}_{\tilde{T}(\epsilon)}\right\}.$ (3.33)
By Lemma 3.4 and 3.5, we can conclude that each iteration $t$ belongs to at least one of the above two sets $\mathcal{F}_{\tilde{T}(\epsilon)}$ and $\mathcal{G}_{\tilde{T}(\epsilon)}$.
Similar to the proof in Section 2.3, we give an upper bound on $\tilde{T}(\epsilon)$ in the following theorem.
Theorem 3.2
Suppose that Assumption 2.1 holds. Let $\{x_{t}\}$ be a sequence generated by the ILMNegCur algorithm with all parameters chosen the same as in Lemma 3.4. We have
$\tilde{T}(\epsilon)$ $\leqslant\tilde{C}_{1}\epsilon^{-3/2}\ln\epsilon^{-1}+\tilde{C}_{2}\epsilon^{-3/2}+1=\tilde{\mathcal{O}}(\ell^{1.5}\rho^{0.5}\mu^{-1.5}\epsilon^{-1.5}),$
$\sum_{t=0}^{\tilde{T}(\epsilon)}N_{t}$ $\leqslant 2\sqrt{\kappa}\ln\frac{\sqrt{\kappa+1}\|y_{-1}-y^{*}(x_{0})\|}{A}+2\sqrt{\kappa}\tilde{T}(\epsilon)\ln\frac{\sqrt{\kappa+1}(A+\kappa r\sqrt{\tilde{G}})}{A},$
where $\tilde{C}_{1}=288\sqrt{L_{2}}(P(x_{0})-P^{*})/\ln 2$, $\tilde{C}_{2}=(\ln\tilde{G}+\ln 2)\tilde{C}_{1}$, $\tilde{G}:=\max\big{\{}\frac{55}{8}\tilde{C},1+\frac{L_{1}}{2\sqrt{L_{2}}}\hat{C}^{1/2}+\tilde{C},2+\frac{L_{1}}{2\sqrt{L_{2}}}\big{\}}$ with $\tilde{C}:=\left(288\sqrt{L_{2}}\left(P(x_{0})-P^{*}\right)\right)^{2/3}$, and $A$ is defined as in Lemma 2.4.
Next, we give the Hessian-vector products required by the ILMNegCur algorithm. Note that, according to Lemmas 9 and 11 in Royer2018ComplexityAO , in each outer iteration, both the randomized Lanczos procedure and the conjugate gradient (CG) procedure require at most $\min\big{\{}n,\tilde{\mathcal{O}}(L_{1}^{0.5}L_{2}^{-0.25}\epsilon^{-0.25})\big{\}}$ Hessian-vector products to obtain a direction $s_{t}$ that satisfies conditions (3.20) and (3.22) with high probability. We immediately obtain the following corollary.
Corollary 3.1
Suppose that Assumption 2.1 holds, and the matrix inverse computation $(\nabla_{yy}^{2}f(x_{t},y_{t}))^{-1}$ in $H_{t}$ is approximated via $\tilde{\mathcal{O}}(\sqrt{\kappa})$ terms of Chebyshev polynomials as in Luo2022FindingSS . Let $\{x_{t}\}$ be a sequence generated by the ILMNegCur algorithm and all other parameters chosen the same as in Lemma 3.4. Then, $x_{\tilde{T}(\epsilon)}$ is an $\mathcal{O}(\epsilon,\sqrt{\epsilon})$-second-order stationary point with high probability guarantee. the total number of Hessian-vector products required is at most $\tilde{\mathcal{O}}(\ell^{2.25}\rho^{0.25}\mu^{-1.75}\epsilon^{-1.75})$.